OpenStudy (anonymous):

What are the x-coordinates of the solutions to this system of equations A. -4 and 0 B. 4 and 0 C. -3 and -5 D. 3 and -5

OpenStudy (anonymous):

X^2+y^2=25 Y=2x-5

OpenStudy (sleepyjess):

Ok, @phi may be able to help you on this

OpenStudy (anonymous):

Ok thanks!

OpenStudy (sleepyjess):

No problem! Welcome to OpenStudy!

OpenStudy (anonymous):

Thanks!

OpenStudy (phi):

you can solve the system, or you can test the choices. to test choice A (for example) start with x=-4 and use the second equation to find y can you do that ?

OpenStudy (anonymous):

I'm not sure what to do

OpenStudy (anonymous):

At allll

OpenStudy (phi):

start with y=2x-5 if x= -4 , what is y ? do find out , erase the "x", and put in -4 (remember 2x means 2 times x) can you do that ?

OpenStudy (anonymous):

Y=-13 ?

OpenStudy (anonymous):

Y=0

OpenStudy (phi):

y=2x-5 with x=-4, we get y= 2* (-4) -5 y= -8 -5 (order of operations means we do multiply first) y= -13 this mean when x=-4 y= -13 now we test if this pair works in the first equation (which is a circle by the way) (-4)^2 + (-13)^2 = 25 I don't need a calculator to know that -13*-13 is bigger than 100 (because 10*10 is 100) and -4*-4 is 16. we will not get 25. all of that means x=-4 is not part of the answer. so choice A is not the answer

OpenStudy (phi):

now let's test choice B let x=4 and find y, using y=2x-5 can you find y ?

OpenStudy (phi):

replace the x with 4 in y=2x-5 and simplify

OpenStudy (anonymous):

Y=3

OpenStudy (phi):

ok. so for choice B, we have x=4, y=3 can you use those numbers in the first equation \[ x^2 + y^2 = 15 \] ?

OpenStudy (anonymous):

No 16+9 =15 it doesn't work

OpenStudy (phi):

oh, I typed in the first equation wrong.

OpenStudy (phi):

can you use the correct equation and try again ?

OpenStudy (phi):

X^2+y^2=25